# Stochastic analysis of finite-temperature effects on cosmological parameters by artificial neural networks

### v1

@@ 3 Numerical analysis
The gradients to optimize the distance function are computed as described in section 2.2.2. Subsequently, we employed the BFGS optimization algorithm to find the minimum of the objective function. Since the explicit form of the distance function is unavailable, a trained ANN was used as a surrogate model to approximate the function and compute the gradients numerically. Accordingly, we implemented the gradient-based optimization iteratively until the algorithm converged and the stopping criterion, given by $||\nabla\mathcal{F}^{L}({\bf x},\widehat{\boldsymbol{\theta}}_{opt})||_{2}<10^{-7}$ is satisfied. Once the ANN-based gradient optimization converges, the estimated optimal coordinates represent the minimum of the approximated distance function. Applying the optimization algorithm to the eight-parameter statistical population, the ANN-based optimization predicted a minimum distance of $y=29.24$ with the corresponding parameter values: $(\Omega_{\Lambda_{2}},\Omega_{\Lambda_{3}},h,\omega_{b},\omega_{cdm},A_{s},n_{s},\tau_{reio})=(-2.069\times 10^{-8},0.00430,0.6752,0.0223,0.1996,2.1\times 10^{-9},0.9659,0.05423).$ By using the same optimization configurations but considering the seven-parameter statistical population, the ANN-based gradient optimization leads to a minimum Planck distance of $y=32.09$. The corresponding parameter values at this minimum were: $(\Omega_{\Lambda_{2}},\Omega_{\Lambda_{3}},h,\omega_{b},\omega_{cdm},A_{s},n_{s},\tau_{reio})=(-2.3356\times 10^{-3},0.00430,0.6768,0.2237,0.1199,2.098\times 10^{-9},0.9657,0.05395).$ By comparing the optimization results under the two statistical populations, we observe that the ANN-based estimate incorporating the finite-T terms attains a lower minimum for the distance function.
Figure 8: Test MSEs of the full and reduced ANN models for predicting the distance function. Each bar represents the mean MSE along with its 95% confidence interval, computed across 50 folds. Reduced models exclude one input parameter at a time to assess its impact on prediction performance.
One can then apply the feature ablation method to evaluate the relative importance of $\Omega_{\Lambda_{2}}$, compared to other established cosmological parameters. Since the explicit form of the distance function is not available, we employed feature ablation using ANNs to estimate its impact in two feature importance numerical studies.
@@
Table 3: MSEs of the full models and reduced models after feature removals using ANN models using 5000 epochs based 700k data points obtained by comping all 8- and 7-parameter statistical populations.
@@
In the next feature ablation analysis, the aim is to evaluate the individual contributions of each input feature to the minimum values of the distance function. Since the objective is to identify the function’s minimum, we focus on a subset of 5000 data points corresponding to the lowest distances from the combined population. We apply the ANNs by using the same architecture as in the previous analysis, but train with 5000 epochs. For uncertainty quantification, we employ 50-fold cross-validation. In each of the 50 runs, 49 folds (4900 data points) are used for training the ANN, and the remaining fold (100 data points) is used for testing. This process is repeated, so that each fold serves once as the test set. In each run, the test MSE for both the full model and reduced models are calculated, each constructed by removing a single feature. This yields 50 MSE values for the full model and for each reduced model. Fig. 8 presents the results of this analysis, where the height of each blue bar shows the mean test MSE and the black lines denote the 95% confidence intervals. The variation in MSE across reduced models reveals that the features contribute unequally to the ANN’s performance in predicting the minimum distance. In addition, removing the feature $\Omega_{\Lambda_{2}}$ increased the average test MSE by approximately 25% compared to the full model, indicating its strong influence. Moreover, when comparing $\Omega_{\Lambda_{2}}$ to some of the existing parameters, we observe that the model with $\Omega_{\Lambda_{2}}$ has a greater impact on model accuracy. This suggests that $\Omega_{\Lambda_{2}}$ is a more important predictor than some of the established cosmological parameters.
@@
With the significance of $\Omega_{\Lambda_{2}}$ confirmed through feature ablation, we now use the simulated annealing approach of section 2.2.4. This method is applied to the statistical population with the quantum gravity feature to determine the minimum of the distance function. The optimization problem involves eight input parameters, with the objective function being unavailable in an explicit analytic form. Instead, it must be approximated stochastically through forward and backward computations of an ANN, as described earlier. Given the integration of multiple numerical methods, it is crucial to account for the sampling variability of parameter estimates and the numerical measurement errors in identifying the minimum. To address this challenge, we employed the stochastic simulated annealing approach within a probabilistic framework.
The optimization process begins with an initial random solution, and perturbations are introduced iteratively using a random walk Metropolis-Hastings algorithm to explore the parameter space. The initial temperature is set to $T_{0}=10^{4}$, and a cooling schedule is applied with a decay rate of $\alpha=0.99$, gradually reducing the acceptance probability for worse solutions as the search progressed.
The ANN predicts the distance response for proposed solutions throughout the optimization, guiding the SA algorithm toward the global minimum. The SA algorithm is run for 20,000 iterations, where 635 transitions are accepted. The results, illustrated in Fig. 9, show the MCMC sequence of accepted minimal distances obtained via ANN-based SA. Ultimately, the algorithm converges to the minimum distance value of $y=28.8111$, with the corresponding optimal coordinates: $(\Omega_{\Lambda_{2}},\Omega_{\Lambda_{3}},h,\omega_{b},\omega_{cdm},A_{s},n_{s},\tau_{reio})=(-2.061\times 10^{-8},0.00425,0.6762,0.0223,0.1196,2.1\times 10^{-9},0.9664,0.056)$. The SA algorithm efficiently explores the complex and noisy domain, with the ANN acting as a surrogate model to approximate the distance function. This ensures robustness in parameter estimation, effectively handling both numerical measurement errors in the method and the sampling variability of the estimated parameters.
@@ 4 Conclusion
We evaluated the effectiveness and robustness of ANN-based methods in estimating cosmological parameters. Our results demonstrate that ANNs effectively capture complex nonlinear dependencies, accurately modeling the cosmological data. Numerical experiments indicate that the introduction of finite-T induced density parameters significantly enhances predictive performance, enabling faster convergence and higher accuracy, even with limited training data. This underscores the potentially significant role of quantum gravitational corrections in refining cosmological predictions. Moreover, our extensive numerical analysis revealed that finite-T corrections can have a greater impact on predictive accuracy than some established cosmological parameters. To further optimize the parameter estimation process, we implemented a simulated annealing algorithm guided by ANN predictions. This stochastic optimization approach efficiently explores the parameter space while accounting for numerical uncertainties in the distance function, yielding reliable estimates for the underlying cosmological parameters.
A central motivation of this work was to examine whether finite-temperature quantum gravity corrections could help mitigate the Hubble tension [17]. Although our results do not resolve the tension outright, they suggest that higher-order finite-temperature contributions may hold the potential for the problem. Further investigation into these effects is therefore warranted. Additionally, a comparison with polarization data - offering complementary constraints on the CMB - will be useful to solidify and refine these conclusions.

### v2

@@ 3 Numerical analysis
The gradients to optimize the distance function are computed as described in section B.2. Subsequently, we employed the BFGS optimization algorithm to find the minimum of the objective function. Since the explicit form of the distance function is unavailable, a trained ANN was used as a surrogate model to approximate the function and compute the gradients numerically. Accordingly, we implemented the gradient-based optimization iteratively until the algorithm converged and the stopping criterion, given by $\nabla\mathcal{F}^{L}({\bf x},\widehat{\boldsymbol{\theta}}_{opt})||_{2}<10^{-7}$ is satisfied. (See Appendix B.) Once the ANN-based gradient optimization converges, the estimated optimal coordinates represent the minimum of the approximated distance function. Applying the optimization algorithm to the eight-parameter statistical population, the ANN-based optimization predicted a minimum distance of $y=29.24$ with the corresponding parameter values: $(\Omega_{\Lambda_{2}},\Omega_{\Lambda_{3}},h,\omega_{b},\omega_{cdm},A_{s},n_{s},\tau_{reio})=(-2.069\times 10^{-8},0.00430,0.6752,0.02240,0.1996,2.1007\times 10^{-9},0.9660,0.05424).$ By using the same optimization configurations but considering the seven-parameter statistical population, the ANN-based gradient optimization leads to a minimum Planck distance of $y=32.09$. The corresponding parameter values at this minimum were:
$(\Omega_{\Lambda_{3}},h,\omega_{b},\omega_{cdm},A_{s},n_{s},\tau_{reio})=(0.0023356,0.6769,0.2237,0.1199,2.0988\times 10^{-9},0.9658,0.05396).$ By comparing the optimization results under the two statistical populations, we observe that the ANN-based estimate of the eight parameter model attains a lower minimum for the distance function. One can then apply the feature ablation method reviewed in section B.3 to evaluate the relative importance of $\Omega_{\Lambda_{2}}$, compared to other established cosmological parameters. Since the explicit form of the distance function is not available, we employed feature ablation using ANNs to estimate its impact in two feature importance numerical studies.
@@
Table 3: This result is based on the entire 700K data. MSEs of the full models and reduced models after feature removals using ANN models using 5000 epochs based 700k data points obtained by combining all 8- and 7-parameter statistical populations.
@@
In the next feature ablation analysis, the aim is to evaluate the individual contributions of each input feature to the minimum values of the distance function. Since the objective is to identify the function’s minimum, we focus on a subset of 5000 data points corresponding to the lowest distances from the combined population. We apply the ANNs by using the same architecture as in the previous analysis, but train with 5000 epochs. For uncertainty quantification, we employ 50-fold cross-validation. In each of the 50 runs, 49 folds (4900 data points) are used for training the ANN, and the remaining fold (100 data points) is used for testing. This process is repeated, so that each fold serves once as the test set. In each run, the test MSE for both the full model and reduced models are calculated, each constructed by removing a single feature. This yields 50 MSE values for the full model and for each reduced model.
Figure 7: This result is based on the first 5000 shortest distances. Test MSEs of the full and reduced ANN models for predicting the distance function. Each bar represents the mean MSE along with its 95% confidence interval, computed across 50 folds. Reduced models exclude one input parameter at a time to assess its impact on prediction performance.
Fig. 7 presents the results of this analysis, where the height of each blue bar shows the mean test MSE and the black lines denote the 95% confidence intervals. The variation in MSE across reduced models reveals that the features contribute unequally to the ANN’s performance in predicting the minimum distance. In addition, removing the feature $\Omega_{\Lambda_{2}}$ increased the average test MSE by approximately 25% compared to the full model, indicating its strong influence. Moreover, when comparing $\Omega_{\Lambda_{2}}$ to some of the parameters, we observe that the model with $\Omega_{\Lambda_{2}}$ has a greater impact on model accuracy.
Figure 7 presents a local sensitivity analysis based on a computationally intensive numerical experiment. Its purpose is to examine the relative importance of each parameter in achieving the minimum distance - essentially identifying which parameters are most critical for finding the ’optimal’ values. This analysis focuses specifically on how individual parameters influence the optimization process near that optimum point. Table 3, by contrast, is based on a global analysis using all 700,000 data points across the full parameter space (combining both the 7-parameter and 8-parameter models). The objective of Table 3 is not optimization, but rather to assess which parameters are most important for accurately estimating the entire Planck curve over its full domain in $\mathbb{R}^{8}$. Due to the large dataset size, the Artificial Neural Network (ANN) was trained only once for Table 3. Parameter importance was then evaluated using a ’leave-one-out’ strategy: each parameter was excluded in turn, and the resulting increase in estimation error was recorded. The magnitude of this error increase serves as a quantitative measure of each parameter’s contribution to the accurate global reconstruction of the Planck curve.
The results shown in Table 3 and Fig. 7 suggest that $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$ play an important role in determining the fine structure of the curve. Table 3 shows that the large-scale shape of the curve is primarily determined by the standard parameters. However, once attention is restricted to curves that closely match the Planck 2018 result, the parameters $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$ become essential.
@@
With the significance of $\Omega_{\Lambda_{2}}$ confirmed through feature ablation, we now use the simulated annealing approach of section B.4. This method is applied to the statistical population with the quantum gravity feature to determine the minimum of the distance function. The optimization problem involves eight input parameters, with the objective function being unavailable in an explicit analytic form. Instead, it must be approximated stochastically through forward and backward computations of an ANN, as described earlier. Given the integration of multiple numerical methods, it is crucial to account for the sampling variability of parameter estimates and the numerical measurement errors in identifying the minimum. To address this challenge, we employed the stochastic simulated annealing approach within a probabilistic framework.
The optimization process begins with an initial random solution, and perturbations are introduced iteratively using a random walk Metropolis-Hastings algorithm to explore the parameter space. The initial temperature is set to $T_{0}=10^{4}$, and a cooling schedule is applied with a decay rate of $\alpha=0.99$, gradually reducing the acceptance probability for worse solutions as the search progressed. The ANN predicts the distance response for proposed solutions throughout the optimization, guiding the SA algorithm toward the global minimum. The SA algorithm is run for 20,000 iterations, where 635 transitions are accepted. The results, illustrated in Fig. 8, show the MCMC sequence of accepted minimal distances obtained via ANN-based SA. Ultimately, the algorithm converges to the minimum distance value of $y=28.8111$, with the corresponding optimal coordinates:
$(\Omega_{\Lambda_{2}},\Omega_{\Lambda_{3}},h,\omega_{b},\omega_{cdm},A_{s},n_{s},\tau_{reio})$ (22)
$=$ $(-2.061\times 10^{-8},0.00425,0.6762,0.0223,0.1196,2.1\times 10^{-9},0.9664,0.056).$
The SA algorithm efficiently explores the complex and noisy domain, with the ANN acting as a surrogate model to approximate the distance function. This ensures robustness in parameter estimation, effectively handling both numerical measurement errors in the method and the sampling variability of the estimated parameters.
@@ 4 Conclusion
We evaluated the effectiveness and robustness of ANN-based methods in estimating cosmological parameters. Our results demonstrate that ANNs are capable of capturing complex nonlinear dependencies and accurately modeling cosmological data. Numerical experiments indicate that the introduction of finite-temperature-induced density parameters enhances predictive performance, enabling faster convergence and improved accuracy, even when training data is limited. This highlights the potential relevance of quantum gravitational corrections in refining cosmological predictions. Notably, our extensive numerical analysis revealed that finite-temperature corrections can have a more substantial impact on predictive accuracy than some well-established cosmological parameters. To further optimize parameter estimation, we incorporated a simulated annealing algorithm guided by ANN predictions. This stochastic optimization strategy efficiently navigates the parameter space while accounting for numerical uncertainties in the distance function, producing robust estimates for the underlying cosmological parameters.
A central motivation of this work was to explore whether finite-temperature quantum-gravity corrections could help alleviate the Hubble tension [25]. While our findings do not resolve the tension outright (see, e.g., [44] for evidence of resolution in dynamical vacuum models), they suggest that higher-order finite-temperature effects may play a meaningful role, warranting further investigation. A key limitation of our present analysis is that, although simulated annealing identifies best-fit parameters, it does not provide likelihood curvature and therefore cannot yield statistical uncertainties; moreover, because the distribution of the distance function is unknown, a classical likelihood cannot be constructed without imposing arbitrary assumptions. [^11] In addition, not all digits reported in the Planck minimum-theory spectrum file COM_PowerSpect_CMB-base-plikHM-TTTEEE-lowl-lowE-lensing-minimum-theory_R3.01.txt are expected to be statistically significant, which limits the meaningful numerical precision of any parameter comparison. To avoid these issues, we reformulated the problem using an artificial neural network (ANN) trained to approximate the distance function across the high-dimensional parameter space in both the 7- and 8-parameter settings, thereby enabling ANN-based numerical optimization to locate the global minimum and the associated cosmological parameter estimates without assuming any particular distribution. To quantify uncertainty and assess the relative importance of each parameter, we implemented a numerical ANN-based feature-ablation procedure, which provides 95% confidence intervals for the prediction MSE and highlights the contribution of each cosmological parameter to the distance function; the resulting intervals are shown in Fig. 8. Looking ahead [45], implementing a full Bayesian inference analysis - using packages such as Cobaya or MontePython - will be a priority for statistically quantifying parameter constraints and validating the ANN-guided results presented here, and a dedicated comparison with polarization data will be essential for further solidifying and refining these conclusions.
@@ Acknowledgments
Appendix A Demonstration of non-degeneracy of $\Omega_{\Lambda_{3}}$ and $\Omega_{K}$
In this appendix, we explicitly demonstrate the non-degeneracy of $\Omega_{\Lambda_{3}}$ and $\Omega_{K}$. Let us choose $(\Omega_{\Lambda_{2}},\Omega_{\Lambda_{3}},h,\omega_{b},\omega_{cdm},A_{s},n_{s},\tau_{reio})$ as follows:
$\hskip 195.12877pt\Omega_{\Lambda_{2}}=0.00000017$
$\hskip-21.68121pth\!=\!0.67810,\;\omega_{b}\!=\!0.02238280,\;\omega_{cdm}\!=\!0.1201075,\;A_{s}\!=\!2.100549\!\times\!10^{-9},\;n_{s}\!=\!0.9660499,\;\tau_{reio}\!=\!0.054308.$
As for $\Omega_{\Lambda_{3}},\Omega_{K}$, let us try the following two combinations:
$\Omega_{\Lambda_{3}}=0.001,\;\Omega_{K}=0$
$\Omega_{\Lambda_{3}}=0,\;\Omega_{K}=0.001.$ (A.2)
Figs. 9 and 10 show the first 10 columns of the cl_lensed.dat files obtained by running corresponding ini files.
Figure 9: $\Omega_{\Lambda_{3}}=0.001,\;\Omega_{K}=0$.
Figure 10: $\Omega_{\Lambda_{3}}=0,\;\Omega_{K}=0.001$.
As these figures show, the entries for the two cases are distinct.
Appendix B Statistical methods
Let us review the statistical and ML techniques employed in the main analysis. In the present study, we employ ANNs to predict the distance between the 2018 Planck data [18] and data generated by the modified CLASS with input cosmological parameter values. This ANN-based model is further used for sensitivity analysis through feature ablation, as well as for probabilistic optimization of cosmological parameters. Fig. 11 represents the schematic diagram of the ANNs with eight nodes in the input layer representing the eight cosmological parameters and four hidden layers, each with 16 nodes and the output layer with one node of the distance. Once the ANN is trained, we use gradient-based optimization to determine the set of covariates that minimize the predicted distance. We then explain the feature ablation strategy to discuss the relative importance of cosmological parameters in the predictions. Finally, we developed the ANN-based probabilistic optimization to find the minimum of the distance, considering the numerical measurement errors.
B.1 Artificial neural networks
Let $\mathcal{F}^{L}(\mathbf{x},\boldsymbol{\theta})$ represent a deep neural network with $L$ layers that transforms an input from $\mathbb{R}^{d_{\text{in}}}$ into an output in $\mathbb{R}^{d_{\text{out}}}$, where $d_{\text{in}}$ and $d_{\text{out}}$ denote the dimensions of the input and output spaces, respectively. Suppose $m_{l}$ represents the number of neurons in the $l$-th hidden layer of the ANN, where $m_{0}=d_{\text{in}}$ and $m_{L}=d_{\text{out}}$ corresponds to the input and output layer sizes, respectively. Each layer $l$ $(l=1,\dots,L-1)$ is associated with a weight matrix $\mathbf{W}^{l}\in\mathbb{R}^{m_{l}\times m_{l-1}}$, whose element $W^{l}_{j,k}$ denotes the weight connecting neuron $j$ in layer $l$ to neuron $k$ in layer l1l-1. In addition, each layer uses a bias vector $\mathbf{b}^{l}\in\mathbb{R}^{m_{l}}$, where the $j$-th element, $b_{j}^{l}$, represents the bias term associated with the $j$-th neuron in layer $l$. Suppose $\boldsymbol{\theta}=\{\mathbf{W}^{1},\dots,\mathbf{W}^{L},\mathbf{b}^{1},\dots,\mathbf{b}^{L}\}$ denotes the set of all unknown parameters of the ANN. Let $\mathbf{M}^{l}$ denote the response from layer $l$ before activation; hence $\mathbf{M}^{l}$ is given by:
$\mathbf{M}^{l}=\mathbf{W}^{l}\cdot\mathcal{F}^{l-1}(\mathbf{x},\boldsymbol{\theta})+\mathbf{b}^{l},\qquad\text{for }l=1,\dots,L-1.$ (B.1)
$\mathcal{F}^{l}(\mathbf{x},\boldsymbol{\theta})$ $=\sigma(\mathbf{M}^{l}),\qquad l=1,\dots,L-1,$ (B.2)
$\mathcal{F}^{L}(\mathbf{x},\boldsymbol{\theta})$ $=\mathbf{M}^{L}.$
$\mathcal{J}(\mathcal{F}^{L}(\mathbf{x},\boldsymbol{\theta}))=\sum_{i=1}^{n}(f(\mathbf{x}_{i})-\mathcal{F}^{L}(\mathbf{x},\boldsymbol{\theta}))^{2},$ (B.3)
$\frac{\partial\mathcal{J}(\mathcal{F}^{L}(\mathbf{x},\boldsymbol{\theta}))}{\partial\mathbf{M}^{l}}=\left(\mathbf{W}^{l+1}\right)^{\top}\cdot\frac{\partial\mathcal{J}(\mathcal{F}^{L}(\mathbf{x},\boldsymbol{\theta}))}{\partial\mathbf{M}^{l+1}}\odot\sigma^{\prime}(\mathbf{M}^{l}),$ (B.4)
$\frac{\partial\mathcal{J}(\mathcal{F}^{L}(\mathbf{x},\boldsymbol{\theta}))}{\partial\mathbf{M}^{L}}=\frac{\partial\mathcal{J}(\mathcal{F}^{L}(\mathbf{x},\boldsymbol{\theta}))}{\partial\mathcal{F}^{L}(\mathbf{x},\boldsymbol{\theta})}\odot\sigma^{\prime}(\mathbf{M}^{L}).$ (B.5)
B.2 Gradient-based feature optimization with ANN
$\nabla\mathcal{F}^{L}({\bf x},\widehat{\boldsymbol{\theta}}_{opt})=\frac{\partial\mathcal{F}^{L}({\bf x},\widehat{\boldsymbol{\theta}}_{opt})}{\partial{\bf x}}.$ (B.6)
${\bf x}^{(t+1)}={\bf x}^{(t)}-\alpha\nabla\mathcal{F}^{L}({\bf x}^{(t)},\widehat{\boldsymbol{\theta}}_{opt})$ (B.7)
${\bf x}^{(t+1)}={\bf x}^{(t)}-{\bf H}_{t}^{-1}\left(\mathcal{F}^{L}({\bf x}^{(t)},\widehat{\boldsymbol{\theta}}_{opt})\right)\nabla\mathcal{F}^{L}({\bf x}^{(t)},\widehat{\boldsymbol{\theta}}_{opt}),$ (B.8)
where ${\bf H}_{t}\left(\mathcal{F}^{L}({\bf x}^{(t)},\widehat{\boldsymbol{\theta}}_{opt})\right)$ represents an approximation of the Hessian matrix, which ensures a more accurate search direction compared to standard gradient descent. The algorithm is repeatedly alternated until the stopping rule of $\nabla\mathcal{F}^{L}({\bf x},\widehat{\boldsymbol{\theta}}_{opt})||_{2}<\epsilon$ is satisfied, where $||\cdot||_{2}$ represents the $\ell_{2}$ norm.
B.3 Feature ablation method
Let $\{({\bf x}_{i},y_{i});i=1,\ldots,n\}$ and $\{({\bf x}_{j,{\text{t}est}},y_{j,{\text{t}est}});j=1,\ldots,m\}$ denote, respectively, the training and test data of sizes $n$ and $m$ from the representing the Planck distance response function $y=f({\bf x})$ where $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$. Because the objective response function is not explicitly available, as described in section B.1, we first predict the objective function using the ANN methods based on the training data. Let $\mathcal{F}^{L}({\bf x},\widehat{\boldsymbol{\theta}}_{opt})$ denote the ANN-based estimate of the objective function $f({\bf x})$. The prediction accuracy performance of the trained ANN estimate in the presence of all the input features is computed based on MSE using test data [47]. Accordingly, the ANN-based MSE is computed by
$\text{MSE}\left(\mathcal{F}^{L}({\bf x},\widehat{\boldsymbol{\theta}}_{opt})\right)=\frac{1}{m}\sum_{j=1}^{m}\left(y_{j,{\text{t}est}}-\mathcal{F}^{L}({\bf x}_{j,{\text{t}est}},\widehat{\boldsymbol{\theta}}_{opt})\right)^{2}.$ (B.9)
Henceforth, the MSE (B.9) is called the full model MSE, showing the prediction error from an ANN-based estimate of the objective function using all the input features.
In the next step, we iteratively remove the individual feature $x_{j},j=1,\ldots,d$ from the input design matrix. Let ${\bf x}_{-j}$ denote the input features when the $j$-th feature $x_{j}$ has been removed from the input data. We then train the ANN estimator of the objective function similar to the full model and find the ANN-based estimate $\widehat{f({\bf x})}=\mathcal{F}^{L}({\bf x}_{-j},\widehat{\boldsymbol{\theta}}_{opt})$. Accordingly, we compute the reduced model MSE in a similar vein to equation (B.9) using the test data in the absence of the $x_{j}$ coordinates. Let $\text{MSE}\left(\mathcal{F}^{L}({\bf x}_{-j},\widehat{\boldsymbol{\theta}}_{opt})\right)$ denote the test MSE of the reduced model using data after removing the $j$-th coordinate for $j=1,\ldots,d$.
We then compute the feature ablation statistic after removing $x_{j}$ by
$\Delta\text{MSE}_{j}=\text{MSE}\left(\mathcal{F}^{L}({\bf x}_{-j},\widehat{\boldsymbol{\theta}}_{opt})\right)-\text{MSE}\left(\mathcal{F}^{L}({\bf x},\widehat{\boldsymbol{\theta}}_{opt})\right),$ (B.10)
B.4 ANN-based simulated annealing
${\bf x}^{*}={\bf x}^{(k)}+\Delta{\bf x},$
The new candidate $\bf x$$^{*}$ is then evaluated through a stochastic accept-reject method to see if $\bf x$$^{*}$ improves the minimum of the objective function or not. Since the objective function $f({\bf x})$ is unknown, the accept reject is applied to the ANN estimate of the objective function $\widehat{f}({\bf x})=\mathcal{F}^{L}({\bf x},\widehat{\boldsymbol{\theta}}_{opt})$ computed by the method described in section B.1. Since the closed form of the ANN estimate is not available, the improvement of the new candidate $\bf x$$^{*}$ should be evaluated numerically. Hence, the cost difference based on ANN is computed by
$\Delta\mathcal{F}({\bf x}^{*};{\bf x}^{(k)})=\mathcal{F}^{L}({\bf x}^{*},\widehat{\boldsymbol{\theta}}_{opt})-\mathcal{F}^{L}({\bf x}^{(k)},\widehat{\boldsymbol{\theta}}_{opt}).$ (B.11)
[…]
$T_{k+1}=\alpha T_{k},$ (B.13)
where $0<\alpha<1$ is the cooling rate. It is easy to see that in the first iteration of the SA, the high values of the temperature parameter enable the algorithm to explore a wider area of the parameter space to search better for a minimum. If the temperature drops too quickly, the algorithm may not be able to search the entire parameter scape and hence may converge to a suboptimal solution. Conversely, if the temperature is cooled too slowly, the algorithm requires many iterations to go from one area to another of the parameter space. Accordingly, the number of iterations increases, and the $\alpha$ rate controls the cooling temperature for balance exploitation and refinement of the minimum in the parameter space. The algorithm is alternated until no significant improvement is observed in the successive iterations or the algorithm reaches a pre-specified maximum number of iterations.